# Masses and radiative decay widths of the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ and their bottom analogs

### v1

@@ I Introduction
In 2003, the BABAR Collaboration first observed the $D_{s0}^{*}(2317)$ through the $D_{s}^{+}\pi^{0}$ invariant mass spectrum and estimated that its mass is near $2.32$ GeV and its width is small BaBar:2003oey . Shortly afterwards the CLEO Collaboration confirmed the existence of $D_{s0}^{*}(2317)$ and found another state $D_{s1}^{\prime}(2460)$ in the $D_{s}^{+}\pi^{0}$ channel CLEO:2003ggt . They were further confirmed later by Belle and BABAR Belle:2003kup ; Belle:2003guh ; BaBar:2004yux ; BaBar:2003cdx ; BaBar:2006eep . Their isospin and spin parity quantum number $I(J^{P})$ are determined as 0($0(0^{+})$ and $0(1^{+})$ ParticleDataGroup:2024cfk , and their masses and widths are ParticleDataGroup:2024cfk
@@
This article is organized as follows. After the introduction, we present the general formalism of the coupled-channel Schrödinger equation for the masses of the compound states in Sec. II, and the involved interactions are given in detail in Sec. III. We analyze the unquenched effect for $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom partners in Sec. IV. Their radiative decays are then studied in Sec. V. A short summary follows in Section VI.
@@ III Detailed interactions
In this section, we provide the quark-model descriptions of the bare $c\bar{s}$/$\bar{b}s$ cores in Subsec. III.1, the effective Lagrangian calculations of the $D^{(*)}K$/$B^{(*)}\bar{K}$ self-interactions in Subsec. III.2 and the couplings between the bare states and the $D^{(*)}K/B^{(*)}\bar{K}$ in Subsec. III.3, respectively. To implement these detailed interactions into the coupled-channel Schrödinger equations, one can then study the unquenched effects of the $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom analogs.
@@ III.1 Quark-model description of the bare cores
III.2 Two-hadron interactions in effective Lagranian
@@
$\begin{split}&V_{\rm{Contact}}=-\frac{3m_{K}}{4f_{K}^{2}},\\ &V_{\rho}=-\frac{3}{2}\beta g_{V}g_{\rho KK}\frac{1}{{\textbf{q}}^{2}+m_{\rho}},\\ &V_{\omega}=-\frac{1}{2}\beta g_{V}g_{\omega KK}\frac{1}{{\bf q}^{2}+m_{\omega}^{2}},\\ &V_{\sigma}=-g_{s}g_{\sigma}\frac{1}{{\bf q}^{2}+m_{\sigma}^{2}},\\ &\mathcal{V}_{\rm Total}=V_{\rm{Contact}}+V_{\rho}+V_{\omega}+V_{\sigma}\,,\end{split}$ (39)
@@ III.3 Interaction between the bare core and hadron-hadron channel
In the non-relativistic limit, $\hat{H}_{I}$ is equivalent to Chen:2016iyi
@@
Using the wave functions to describe the bare hadrons in previous Subsec. III.1, we can obtain the transition amplitudes of bare cores to hadron-hadron channels in Eq. (LABEL:eq:TranM) with the quark-pair-creation model.
@@ V.2 The radiative decay for pure molecular state
With the above preparations, we can calculate the radiative decay widths of the pure molecular states and the results are shown in Fig. 6. By taking the cutoff $\Lambda^{\prime}$ in the range 0.4$\sim$2.0 GeV, the decay width of $[DK]\to D^{*}_{s}\gamma$ is about 4$\sim$45 keV, and those of $[D^{*}K]\to D^{*}_{s}\gamma$ and $[D^{*}K]\to D_{s}\gamma$ are 0.4$\sim$2.1 keV and 1$\sim$16 keV, respectively. For the bottom sector, the decay widths of $[B\bar{K}]\to B_{s}^{*}\gamma$, $[B^{*}\bar{K}]\to B_{s}^{*}\gamma$, and $[B^{*}\bar{K}]\to B_{s}\gamma$ are about 8$\sim$238 keV, 0.2 keVsimilar-$\sim$1.4 keV, and 4$\sim$250 keV, correspondingly. The ratio between $[DK]\to D_{s}^{*}\gamma$ and $[BK]\to B_{s}^{*}\gamma$ is $\frac{[BK]\to B_{s}^{*}\gamma}{[DK]\to D_{s}^{*}\gamma}\approx 2\sim 3$, which is close to the prediction of Ref.  Faessler:2008vc .
@@ V.3 The radiative decay for the physical state
Experimental data on these processes are still lacking, and different theoretical predictions with parity doubling model Bardeen:2003kt , effective Lagrangian approach Faessler:2007gv ; Faessler:2008vc ; Fu:2021wde , light-cone sum rules Colangelo:2005hv ; Wang:2008wz ; Wang:2006mf , and so on, vary in a very wide range. Ref. Wang:2019ehs treated $B_{s1}^{\prime}$ state as a molecular state and provided the widths of $B_{s1}^{\prime}\to B_{s}^{*}\gamma$ and $B_{s1}^{\prime}\to B_{s}\gamma$ as 0.4$\sim$2.6 and 45.2$\sim$79.8 keV, respectively, which agree relatively well with our results in Case \@slowromancapi@.

### v2

@@ I Introduction
In 2003, the BABAR Collaboration first observed the $D_{s0}^{*}(2317)$ through the $D_{s}^{+}\pi^{0}$ invariant mass spectrum and estimated that its mass is near $2.32$ GeV and its width is small BaBar:2003oey . Shortly afterward the CLEO Collaboration confirmed the existence of $D_{s0}^{*}(2317)$ and found another state $D_{s1}^{\prime}(2460)$ in the $D_{s}^{+}\pi^{0}$ channel CLEO:2003ggt . They were further confirmed later by Belle and BABAR Belle:2003kup ; Belle:2003guh ; BaBar:2004yux ; BaBar:2003cdx ; BaBar:2006eep . Their isospin and spin parity quantum number $I(J^{P})$ are determined as 0($0(0^{+})$ and $0(1^{+})$ ParticleDataGroup:2024cfk , and their masses and widths are ParticleDataGroup:2024cfk
@@
This article is organized as follows. After the introduction, we present the general formalism of the coupled-channel Schrödinger equation for the masses of the compound states in Sec. II, and the involved interactions are given in detail in Sec. III. We analyze the unquenched effect for $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom partners in Sec. IV. Their radiative decays are then studied in Sec. V. A short summary follows in Sec. VI.
@@ III Detailed interactions
In this section, we provide the quark-model descriptions of the bare $c\bar{s}$/$\bar{b}s$ cores in Sec. III.1, the effective Lagrangian calculations of the $D^{(*)}K$/$B^{(*)}\bar{K}$ self-interactions in Sec. III.2 and the couplings between the bare states and the $D^{(*)}K/B^{(*)}\bar{K}$ in Sec. III.3, respectively. To implement these detailed interactions into the coupled-channel Schrödinger equations, one can then study the unquenched effects of the $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom analogs.
@@ III.1 Quark-model description of the bare cores
III.2 Two-hadron interactions in effective Lagrangian
@@
$\begin{split}&V_{\rm{Contact}}=-\frac{m_{K}}{2f_{K}^{2}},\\ &V_{\rho}=-\frac{3}{2}\beta g_{V}g_{\rho KK}\frac{1}{{\textbf{q}}^{2}+m_{\rho}},\\ &V_{\omega}=-\frac{1}{2}\beta g_{V}g_{\omega KK}\frac{1}{{\bf q}^{2}+m_{\omega}^{2}},\\ &V_{\sigma}=-g_{s}g_{\sigma}\frac{1}{{\bf q}^{2}+m_{\sigma}^{2}},\\ &\mathcal{V}_{\rm Total}=V_{\rm{Contact}}+V_{\rho}+V_{\omega}+V_{\sigma}\,,\end{split}$ (39)
@@ III.3 Interaction between the bare core and hadron-hadron channel
In the nonrelativistic limit, $\hat{H}_{I}$ is equivalent to Chen:2016iyi
@@
Using the wave functions to describe the bare hadrons in previous Sec. III.1, we can obtain the transition amplitudes of bare cores to hadron-hadron channels in Eq. (LABEL:eq:TranM) with the quark-pair-creation model.
@@ V.2 The radiative decay for pure molecular state
With the above preparations, we can calculate the radiative decay widths of the pure molecular states and the results are shown in Fig. 6. By taking the cutoff $\Lambda^{\prime}$ in the range 0.4--2.0 GeV, the decay width of $[DK]\to D^{*}_{s}\gamma$ is about 4--45 keV, and those of $[D^{*}K]\to D^{*}_{s}\gamma$ and $[D^{*}K]\to D_{s}\gamma$ are 0.4--2.1 keV and 1--16 keV, respectively. For the bottom sector, the decay widths of $[B\bar{K}]\to B_{s}^{*}\gamma$, $[B^{*}\bar{K}]\to B_{s}^{*}\gamma$, and $[B^{*}\bar{K}]\to B_{s}\gamma$ are about 8--238 keV, 0.2 keV--1.4 keV, and 4--250 keV, correspondingly. The ratio between $[DK]\to D_{s}^{*}\gamma$ and $[BK]\to B_{s}^{*}\gamma$ is $\frac{[BK]\to B_{s}^{*}\gamma}{[DK]\to D_{s}^{*}\gamma}\approx 2\sim 3$, which is close to the prediction of Ref.  Faessler:2008vc .
@@ V.3 The radiative decay for the physical state
Experimental data on these processes are still lacking, and different theoretical predictions with parity doubling model Bardeen:2003kt , effective Lagrangian approach Faessler:2007gv ; Faessler:2008vc ; Fu:2021wde , light-cone sum rules Colangelo:2005hv ; Wang:2008wz ; Wang:2006mf , and so on, vary in a very wide range. Ref. Wang:2019ehs treated $B_{s1}^{\prime}$ state as a molecular state and provided the widths of $B_{s1}^{\prime}\to B_{s}^{*}\gamma$ and $B_{s1}^{\prime}\to B_{s}\gamma$ as 0.4--2.6 and 45.2--79.8 keV, respectively, which agree relatively well with our results in case \@slowromancapi@.